# S2R-HDR: A Large-Scale Rendered Dataset for HDR Fusion

### v1

@@ 6 Conclusion
This paper introduces the S2R-HDR dataset, a large-scale, high-quality resource for HDR reconstruction in dynamic scenes. By providing diverse, controllable, and high-fidelity synthetic data, the dataset addresses the limitations of existing HDR datasets. Additionally, we propose the S2R-Adapter, a novel domain adaptation method that effectively bridges the gap between synthetic and real data, enabling efficient knowledge transfer. Experimental results on both labeled and unlabeled datasets demonstrate that our S2R-HDR dataset and S2R-Adapter significantly enhance the performance of HDR reconstruction models in real-world scenarios. This provides a viable solution for the HDR field, where data acquisition is often limited. Future work will focus on expanding the S2R-HDR dataset to support a wider range of application scenarios.
Supplementary Material
@@ A.1 Experiments Details
When training these methods on our S2R-HDR dataset, we first generate three different exposure LDR images from the original HDR images. Following this, we apply the same data augmentation techniques, and training schedules across all models. Additionally, we introduce random Gaussian noise with $\sigma\in[0.0001,0.001]$ to the lowest exposure image and $\sigma\in[0.00001,0.0001]$ to the middle exposure image.
For the SCTNet [41] architecture, which is based on the Transformer framework, we employ a linear layer as the projection layer of the S2R-Adapter (as illustrated in the left part of Fig. 5) and integrate it into SCTNet’s WindowAttention Linear Layer. In contrast, for the SAFNet [21] architecture, which is based on CNNs, we utilize a 1$\times$1 convolutional layer as the S2R-Adapter’s projection layer and inject it into the network at layers indexed by [3:25:2] and [42:58:4]. For both CNN and Transformer architecture, we set the rank of the shared branch adapter $r_{s}$ to be $1$, and the rank of the transfer branch $r_{t}$ to be $64$, following [24].
In our test-time adaptation experiments, each test sample is processed only once. To assess sample uncertainty as a measure of domain shift, we employ test-time augmentation techniques. Specifically, we augment test samples using a variety of exposure levels: $[-0.1,-0.5,0,0.5,1]$. Additionally, we apply random transformations, including flips, white balance adjustments, and random Gaussian noise. For augmentations involving exposure and white balance, we apply the parameters to the input images following inverse tone mapping. Correspondingly, the inverse transformations are directly applied to the model outputs.
@@ A.2 Analysis of Domain Gap and Adapter
To gain deeper insights into the domain gap between real and rendered data and to better understand what domain adaptation learns, we compute difference maps for models trained on the rendered dataset (S2R-HDR) before and after domain adaptation (S2R-Adapter) to the SCT dataset. As shown in Fig. A1, the differences are primarily concentrated in regions containing trees, grass, and people, while ground, sky, and buildings remain largely unchanged. This suggests that the key discrepancies between real and rendered data mainly arise in texture-rich areas such as human figures and vegetation. The results further confirm that domain adaptation effectively mitigates the domain gap.
@@
Figure A1: Difference maps of models trained on the rendered dataset (S2R-HDR) before and after domain adaptation (S2R-Adapter) to the SCT dataset. The differences are primarily concentrated in texture-rich regions such as trees, grass, and people, while ground, sky, and buildings remain largely unchanged. This highlights that the key domain discrepancies lie in fine textures and demonstrates the effectiveness of domain adaptation in bridging the domain gap.
A.3 Visualization of Effectiveness of S2R-HDR
Figure A2: Visualization results of our S2R-HDR dataset comparison experiments. Models trained on our S2R-HDR dataset exhibit significantly fewer artifacts compared to those trained on the SCT dataset [41] or Challenge123 dataset [21].
We further show the visualization results of our S2R-HDR dataset comparison experiments in Fig. A2, models trained on our S2R-HDR dataset achieve optimal visual quality compared to those trained on other datasets. Additionally, as depicted in the left image of Fig. A2, our dataset effectively mitigates motion occlusion challenges. Similarly, as shown in the right image of Fig. A2, our dataset effectively addresses challenges related to high light fusion.
@@ A.4 Additional Results on Real-Capture Images
Figure A3: Visualization results on real-captured data without ground truth. Our approach effectively reduces artifacts in highlight areas and alleviates ghosting in nighttime scenarios.
We further provide a visual comparison using real-captured data without ground truth in Fig. A3. Our approach effectively reduces artifacts in challenging scenarios.
A.5 Additional Data Effectiveness Comparison Experiments
To validate the effectiveness of our S2R-HDR dataset, we used SCTNet [41] as the baseline model and conducted experiments on two datasets: the earliest dataset from Kalantari [19] and the Real-HDRV [37] dataset, which, although the largest, is less commonly used. The results, as shown in Tab. A1, with simple fine-tuning, our dataset consistently delivers the best results.
Table A1: Experimental results of data effectiveness comparison on the Kalantari [19] and Real-HDRV [37] Datasets.
@@ A.5 Additional Data Effectiveness Comparison Experiments
PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$ PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$
Train on Kalantari 35.03 44.75 0.9615 0.9941 35.31 25.79 0.9672 0.9621
Train on Real-HDRV 35.37 46.13 0.9651 0.9949 36.41 26.42 0.9711 0.9674
@@
A.6 Additional Dataset Scale Comparison
@@
We include a comparison of dataset scales between our S2R-HDR dataset and the datasets from SCT [41], Challenge123 [21], and Kalantari [19], as presented in Table A2.
@@
Table A2: Comparison of dataset scale. We compare our S2R-HDR dataset with SCT [41], Challenge123 [21], and Kalantari [19].
SCT [41] Challenge123 [21] Kalantari [19] S2R-HDR
Dataset size 144 123 89 24,000
Table A3: Metrics to assess the diversity of different HDR datasets.
FHLP Fraction of HighLight Pixel [12]
EHL Extent of HighLight [12]
SI Spatial Information [36]
CF ColorFulness [13]
stdL standard deviation of Luminance [12]
ALL Average Luminance Level [12]
DR Dynamic Range: the log10 differences between the highest 2% luminance and the lowest 2% luminance. [18]

### v2

@@ 6 Conclusion
This paper introduces the S2R-HDR dataset, a large-scale, high-quality resource for HDR fusion in dynamic scenes. By providing diverse, controllable, and high-fidelity synthetic data, the dataset addresses the limitations of existing HDR datasets. Additionally, we propose the S2R-Adapter, a novel domain adaptation method that effectively bridges the gap between synthetic and real data, enabling efficient knowledge transfer. Experimental results on both labeled and unlabeled datasets demonstrate that our S2R-HDR dataset and S2R-Adapter significantly enhance the performance of HDR fusion models in real-world scenarios. This provides a viable solution for the HDR field, where data acquisition is often limited. Future work will focus on expanding the S2R-HDR dataset to support a wider range of application scenarios.
7 Reproducibility statement
To ensure the reproducibility of our findings, detailed implementation instructions are provided in Section 5 and Appendix A.1. In addition, the source code and datasets are publicly available at URL: https://openimaginglab.github.io/S2R-HDR/.
8 Acknowledgement
This work is supported/funded in part by the Shanghai Artificial Intelligence Laboratory, the Centre for Perceptual and Interactive Intelligence (CPII) Ltd., a CUHK-led InnoCentre under the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government, RGC Early Career Scheme (ECS) No. 24209224, CUHK-CUHK(SZ)-GDSTC Joint Collaboration Fund No. 2025A0505000053. We thank Yixuan Li for helpful discussions on Unreal Engine rendering. We also thank Haoran Yang for his assistance in capturing the real-world dataset.
@@ A.1 Experiments Details
When training these methods on our S2R-HDR dataset, we first generate three different exposure ({-2, 0, +2)}, {-3, 0, +3}) LDR images from the original HDR images. Following this, we apply the same data augmentation techniques, and training schedules across all models. Additionally, we introduce random Gaussian noise with $\sigma\in[0.0001,0.001]$ to the lowest exposure image and $\sigma\in[0.00001,0.0001]$ to the middle exposure image.
For the SCTNet (Tel et al., 2023) architecture, which is based on the Transformer framework, we employ a linear layer as the projection layer of the S2R-Adapter (as illustrated in the left part of Figure 4) and integrate it into SCTNet’s WindowAttention Linear Layer. In contrast, for the SAFNet (Kong et al., 2024) architecture, which is based on CNNs, we utilize a $\times1$ convolutional layer as the S2R-Adapter’s projection layer and inject it into the network at layers indexed by [3:25:2] and [42:58:4]. For both CNN and Transformer architecture, we set the rank of the shared branch adapter $r_{s}$ to be $1$, and the rank of the transfer branch $r_{t}$ to be $64$, following Liu et al. (2023a).
In our test-time adaptation experiments, each test sample is processed only once. To assess sample uncertainty as a measure of domain shift, we employ test-time augmentation techniques. Specifically, we augment test samples using a variety of exposure levels: [0.1,0.5,0,0.5,1][-0.1,-0.5,0,0.5,1]. Additionally, we apply random transformations, including flips, white balance adjustments, and random Gaussian noise. For augmentations involving exposure and white balance, we apply the parameters to the input images following inverse tone mapping. Correspondingly, the inverse transformations are directly applied to the model outputs. This results in $N$ augmented outputs per sample, from which we compute the variance across these outputs to quantify uncertainty $\mathcal{U}(x)$.
@@ A.2 Analysis of Domain Gap and Adapter
Figure 8: The distribution of our S2R-HDR dataset and real captured HDR datasets (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024). Following the approach outlined in Shu et al. (2024); Guo et al. (2023); Hu et al. (2022b), for each HDR image, we first extract 7-dimensional features listed in Table 18 that capture key aspects of HDR, including the extent of dynamic range, intra-frame diversity, and the overall style of the HDR images. These features are then projected into a 2D space using t-SNE (Van der Maaten and Hinton, 2008) for visualization. Regarding the concern about real-dataset sample size, we analyzed all real images (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024) without sampling bias.
Distribution of our S2R-HDR dataset vs. real datasets. To better understand the distribution of our S2R-HDR dataset compared to existing real-world HDR datasets (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024), we extract seven-dimensional features representing key HDR characteristics, including dynamic range, intra-frame diversity, and overall image style. These features are visualized using t-SNE (Van der Maaten and Hinton, 2008), as illustrated in Figure 8. The results demonstrate that our S2R-HDR dataset spans a significantly broader range of HDR scenarios, reflecting richer variations in lighting conditions, scene dynamics, and environmental styles.
Why we need an adapter. Despite the broader coverage of S2R-HDR, there remains a noticeable domain gap between our synthetic data and real HDR datasets, as highlighted in Figure 8. This gap primarily stems from differences in texture representation, environmental details, and natural lighting conditions. To address this, we introduce an S2R-Adapter to bridge the disparity, enabling models trained on synthetic data to generalize effectively to real-world scenes.
@@
What the adapter learns. To gain deeper insights into the domain gap between real and rendered data and to better understand what domain adaptation learns, we compute difference maps for models trained on the rendered dataset (S2R-HDR) before and after domain adaptation (S2R-Adapter) to the SCT dataset. As shown in Figure 9, the differences are primarily concentrated in regions containing trees, grass, and people, while ground, sky, and buildings remain largely unchanged. This suggests that the key discrepancies between real and rendered data mainly arise in texture-rich areas such as human figures and vegetation. The results further confirm that domain adaptation effectively mitigates the domain gap.
@@
Figure 9: Difference maps of models trained on the rendered dataset (S2R-HDR) before and after domain adaptation (S2R-Adapter) to the SCT dataset. The differences are primarily concentrated in texture-rich regions such as trees, grass, and people, while ground, sky, and buildings remain largely unchanged. This highlights that the key domain discrepancies lie in fine textures and demonstrates the effectiveness of domain adaptation in bridging the domain gap.
A.3 Domain Adaptation Method Comparison (Fine-tune VS S2R-Adapter)
As emphasized, the key challenge lies in addressing the domain gap. We summarize the domain adaptation experiments using fine-tuning and the S2R-Adapter. As shown in Table 8, regardless of the method employed, mitigating the domain gap consistently leads to notable performance improvements.
As shown in Table 8, the performance of SCTNet and SAFNet was evaluated on the SCT and Challenge123 datasets after being trained on different datasets. Since SCTNet is specifically designed for the SCT dataset, it already performs well on this dataset. Consequently, using the S2R-Adapter results in only a modest improvement compared to fine-tuning SCTNet on the SCT dataset. However, on the Challenge123 dataset, the use of the S2R-Adapter leads to a more substantial improvement over fine-tuning SCTNet alone, with PSNR-$\mu$ increasing from 42.1 dB to 42.58 dB. Similarly, SAFNet, which is designed for the Challenge123 dataset, shows a more significant performance improvement on the SCT dataset, with PSNR-$\mu$ increasing from 43.03 dB to 43.33 dB. The key challenge remains addressing the domain gap, and mitigating this gap consistently results in significant performance gains.
We further conduct comparative experiments to evaluate the effectiveness of our dataset against existing datasets in an adapter-based setting. Specifically, we first pre-train the SAFNet (Kong et al., 2024) on the S2R-HDR and Challenge123 (Kong et al., 2024) datasets, respectively, and then adapt them to the SCT dataset using the S2R-Adapter. As shown in Table 9, models pretrained on the S2R-HDR dataset achieve better performance after adaptation. Moreover, we observe that using the S2R-Adapter consistently outperforms both training solely on the SCT dataset and joint training on SCT and Challenge123 datasets, demonstrating the advantage of the proposed adaptation strategy.
Moreover, as shown in Table 7 in the main text, our S2R-Adapter not only improves performance but also offers better knowledge control, with reduced knowledge forgetting compared to fine-tuning. In addition to fine-tuning on labeled real data, our S2R-Adapter demonstrates superior performance in test-time adaptation, allowing flexible adaptation to unseen data without the need for ground-truth labels. This is particularly beneficial for HDR fusion tasks, where large-scale, accurately labeled real-world datasets are often unavailable. By pretraining on large synthetic datasets and adapting to any unlabeled target distribution during test time, our approach exhibits greater flexibility and generalization.
Table 8: Experimental results on the effectiveness of the S2R-HDR dataset. We test the cross-dataset generalization of models trained on different datasets. Models trained on our S2R-HDR dataset achieve superior generalization and require only minimal fine-tuning on SCT or Challenge123 to reach state-of-the-art performance. The best results are in bold.
Methods Training Testing on SCT (Tel et al., 2023) Testing on Challenge123 (Kong et al., 2024)
$\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
SCTNet (Tel et al., 2023) SCT (Tel et al., 2023) 42.55 47.51 0.9850 0.9952 34.57 25.07 0.9753 0.9599
Challenge123 (Kong et al., 2024) 32.60 35.93 0.9535 0.9639 40.65 28.73 0.9882 0.9721
SCT & Challenge123 40.00 42.79 0.9800 0.9935 40.04 28.21 0.9898 0.9750
S2R-HDR 34.83 42.32 0.9526 0.9933 41.49 30.37 0.9862 0.9796
S2R-HDR Fine-tune on SCT and Challenge123 41.40 46.37 0.9820 0.9960 41.93 30.33 0.9907 0.9796
S2R-HDR Fine-tune on SCT or Challenge123 43.22 47.28 0.9872 0.9961 42.10 30.18 0.9914 0.9798
S2R-HDR w S2R-Adapter 43.24 48.32 0.9872 0.9962 42.58 30.68 0.9915 0.9805
SAFNet (Kong et al., 2024) SCT (Tel et al., 2023) 42.66 48.38 0.9831 0.9955 34.26 25.50 0.9718 0.9590
Challenge123 (Kong et al., 2024) 35.14 38.77 0.9619 0.9868 41.88 29.73 0.9897 0.9784
SCT & Challenge123 42.12 45.14 0.9853 0.9941 41.61 29.72 0.9901 0.9788
S2R-HDR 34.89 43.85 0.9500 0.9939 42.75 32.11 0.9872 0.9822
S2R-HDR Fine-tune on SCT and Challenge123 42.97 47.84 0.9861 0.9960 42.91 31.05 0.9906 0.9805
S2R-HDR Fine-tune on SCT or Challenge123 43.03 48.79 0.9831 0.9958 43.30 31.59 0.9914 0.9819
S2R-HDR w S2R-Adapter 43.33 48.90 0.9864 0.9959 43.43 31.84 0.9915 0.9824
Table 9: Effectiveness of S2R-HDR and S2R-Adapter. Pretraining SAFNet (Kong et al., 2024) on S2R-HDR and adapting to SCT (Tel et al., 2023) yields the best overall results.
Training $\mu PNSR$-$\ell$ $\mu SSIM$-$\ell$
SCT (Tel et al., 2023) 42.66 48.38 0.9831 0.9955
Challeng123 (Kong et al., 2024) 35.14 38.77 0.9619 0.9868
SCT & Challenge 42.12 45.14 0.9853 0.9941
S2R-HDR with S2R-Adapter on SCT 43.44 48.90 0.9864 0.9959
Challenge123 with S2R-Adapter on SCT 42.92 47.05 0.9856 0.9951
@@ A.4 Additional Results on Real-Capture Images
Figure 10: Visualization results on real-captured data without ground truth. Our approach effectively reduces artifacts in highlight areas and alleviates ghosting in nighttime scenarios.
We further provide a visual comparison using real-captured data without ground truth in Figure 10. Our approach effectively reduces artifacts in challenging scenarios.
A.5 Visualization of Effectiveness of S2R-HDR
Figure 11: Visualization results of our S2R-HDR dataset comparison experiments. Models trained on our S2R-HDR dataset exhibit significantly fewer artifacts compared to those trained on the SCT dataset (Tel et al., 2023) or Challenge123 dataset (Kong et al., 2024).
We further show the visualization results of our S2R-HDR dataset comparison experiments in Figure 11, models trained on our S2R-HDR dataset achieve optimal visual quality compared to those trained on other datasets. Additionally, as depicted in the left image of Figure 11, our dataset effectively mitigates motion occlusion challenges. Similarly, as shown in the right image of Figure 11, our dataset effectively addresses challenges related to high light fusion.
A.6 Additional Data Effectiveness Comparison Experiments
To validate the effectiveness of our S2R-HDR dataset, we used SCTNet (Tel et al., 2023) as the baseline model and conducted experiments on the Real-HDRV (Deghosting) (Shu et al., 2024) dataset, which, although the largest, is less commonly used. The results, as shown in Table 10, with simple fine-tuning, our dataset consistently delivers the best results.
Table 10: Experimental results of data effectiveness comparison with the Real-HDRV (Deghosting) (Shu et al., 2024) Datasets.
@@ A.6 Additional Data Effectiveness Comparison Experiments
$\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
Train on Real-HDRV (Deghosting) 35.37 46.13 0.9651 0.9949 36.41 26.42 0.9711 0.9674
@@
We also use SCTNet (Tel et al., 2023) as a baseline model and conduct experiments on the earliest Kalantari (Kalantari and Ramamoorthi, 2017) dataset. SCTNet is retrained on the entire dataset, and the results, as shown in Table 11, demonstrate that with simple fine-tuning, our dataset consistently yields the best performance.
Table 11: Experimental results of data effectiveness comparison on the earliest Kalantari (Kalantari and Ramamoorthi, 2017) Datasets.
Methods Training $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
SCTNet Kalantari 44.13 42.32 0.9916 0.9890
S2R-HDR 41.41 36.68 0.9859 0.9787
S2R-HDR Fine-tune on Kalantari 44.32 43.16 0.9923 0.9911
A.7 Additional Experiments on the Effectiveness of S2R-Adapter
We conducted additional experiments on more datasets to further evaluate the effectiveness of S2R-Adapter. These experiments include test-time adaptation without ground truth on Kalantari’s (Kalantari and Ramamoorthi, 2017), Sen’s (Sen et al., 2012), and Tursun’s (Tursun et al., 2016) datasets.
We first performed test-time adaptation without ground-truth labels on Kalantari’s dataset using the SAFNet. Table 12 presents the results.
Table 12: Test-time adaptation results on Kalantari (Kalantari and Ramamoorthi, 2017) dataset.
Methods $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
SAFNet trained on SCT 34.75 39.41 0.9806 0.9821
SAFNet trained on Challenge123 40.31 37.19 0.9843 0.9772
SAFNet trained on S2R-HDR 43.06 40.63 0.9887 0.9878
SAFNet trained on S2R-HDR with S2R-Adapter 43.29 41.488 0.9889 0.9890
We also conducted test-time adaptation on Sen’s (Sen et al., 2012) and Tursun’s (Tursun et al., 2016) datasets. We report the non-reference IQA metric MUSIQ (Ke et al., 2021) scores in Table 13.
Table 13: Test-time adaptation results (MUSIQ scores) on Sen’s (Sen et al., 2012) and Tursun’s (Tursun et al., 2016) datasets.
Methods MUSIQ on Tursun’s MUSIQ on Sen’s
SAFNet trained on SCT 63.94 65.71
SAFNet trained on Challenge123 63.42 65.65
SAFNet trained on S2R-HDR 63.32 65.97
SAFNet trained on S2R-HDR with S2R-Adapter 65.30 67.45
A.8 Additional Results on HDR Video Reconstruction Task
We also conduct experiments on HDR video reconstruction task to validate the effectiveness of our S2R-HDR dataset. We use recent open-sourced methods, HDRFlow (Xu et al., 2024), as our baseline model, and conduct training experiments on Vimeo-90K (Xue et al., 2019), Sintel (Butler et al., 2012), and Real-HDRV (Shu et al., 2024) datasets, testing on DeepHDRVideo (Chen et al., 2021) dataset. The results, as shown in Table 14, with simple fine-tuning, our dataset consistently delivers the best results.
Table 14: Experimental results of data effectiveness comparison on the HDR video reconstruction task. The HDRFlow method is trained on different datasets and tested on the DeepHDRVideo (Chen et al., 2021) dataset.
Methods Training $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
HDRFlow (Xu et al., 2024) Vimeo+Sintel 43.25 53.05 0.9520 0.9956
Real-HDRV 43.37 53.05 0.9540 0.9958
S2R-HDR 43.13 52.50 0.9508 0.9953
S2R-HDR Fine-tune on Real-HDRV 43.51 53.52 0.9546 0.9960
Additionally, in Table 15, we tested temporal consistency using the TOG HDR dynamic dataset (Kalantari et al., 2013) with method (Lai et al., 2018), this metric is calculated using Equation 4, which measures the temporal consistency of a video by quantifying the flow warping error between adjacent frames. In this equation, $V_{t}$ represents the original frame, $\hat{V}_{t+1}$ represents the warped frame using optical flow and $M$ is a non-occlusion mask indicating non-occluded regions. The results show that our dataset achieves the best performance in temporal consistency.
Table 15: Experimental results of temporal consistency comparison on the HDR video reconstruction task.
Training Datasets Sintel+Vimeo HDRV S2R-HDR
Temporal Stability Score↓ 0.2201 0.2183 0.1773
$\begin{array}[]{l}E_{\text{warp }}\left(V_{t},V_{t+1}\right)=\frac{1}{\sum_{i=1}^{N}M_{t}^{(i)}}\sum_{i=1}^{N}M_{t}^{(i)}\left\|V_{t}^{(i)}-\hat{V}_{t+1}^{(i)}\right\|_{2}^{2}\\ E_{\text{warp }}(V)=\frac{1}{T-1}\sum_{t=1}^{T-1}E_{\text{warp }}\left(V_{t},V_{t+1}\right)\end{array}$ (4)
A.9 Detailed Dataset Generation Pipeline
We provide a complete and transparent overview of our dataset-generation pipeline in Figure12. Because this pipeline allows us to directly control the lighting parameters of the sources in the UE5 simulation environment (including their positions, intensities, etc.), we can synthesize a highly diverse set of illumination conditions. In addition, we can manipulate the pose and motion of objects in the scene by binding different animation sequences to them, thereby producing a wide variety of motion patterns. With the exception of the use of several high-quality UE5 scene assets that we purchased for scene construction, the rest of the pipeline is fully reproducible.
Figure 12: Distribution of optical flow directions across S2R-HDR.
A.10 Detailed Dataset Scale Comparison
@@
We include a comparison of dataset scales between our S2R-HDR dataset and the datasets from SCT (Tel et al., 2023), Challenge123 (Kong et al., 2024), Kalantari (Kalantari and Ramamoorthi, 2017), and Real-HDRV(Deghosting) (Shu et al., 2024), as presented in Table 16.
@@
Table 16: Comparison of dataset scale. We compare our S2R-HDR dataset with SCT (Tel et al., 2023), Challenge123 (Kong et al., 2024), Kalantari (Kalantari and Ramamoorthi, 2017), and Real-HDRV(Deghosting) (Shu et al., 2024).
SCT (Tel et al., 2023) Challenge123 (Kong et al., 2024) Kalantari (Kalantari and Ramamoorthi, 2017) Real-HDRV(Deghosting) (Shu et al., 2024) S2R-HDR
Dataset size 144 (108/36) 123 (96/27) 89 (74/15) 500 (450/50) 24,000
A.11 Further scene-level Analysis of S2R-HDR
We provide additional scene-level analysis of our proposed dataset S2R-HDR. First, following the same methodology as in (Shu et al., 2024), we analyze the directions of motion using the optical flow. Figure13 reports the proportion of pixels corresponding to different motion directions in S2R-HDR, from which it can be observed that the motion patterns in our dataset are highly diverse. Second, Table19 summarizes the distribution of illumination conditions and locations. In terms of scene semantics, our dataset covers a wide variety of environments, including home interiors, urban streets, churches, rural areas, parks, subway stations, workshops, and more. Regarding the foreground assets, Figure14 further illustrates the richness of our foreground content together with the associated motion patterns.
Figure 13: Distribution of optical flow directions across S2R-HDR.
A.12 Test-time Adaptation Runtime
Like most test-time adaptation (TTA) methods (Kundu et al., 2020; Liu et al., 2023a; Wang et al., 2022; Sun et al., 2020; Wang et al., 2020), our approach incurs additional computational cost due to model weight updates during inference to adapt to the target domain without ground-truth labels. However, this cost is modest and justified by the performance gains.
To quantify the overhead, we evaluated test-time adaptation runtime on SCT test images (resolution 1500×1000) using SAFNet as the base model. Experiments were conducted on a server with an AMD EPYC 7402 (48C) @ 2.8 GHz CPU, 8×NVIDIA RTX 4090 GPUs, 512 GB RAM, running CentOS 7.9. The results are as follows:
Table 17: Test-time Adaptation Runtime
Standard Testing Test-Time Adaptation
Runtime (seconds) 0.3813 0.6128
PSNR-tested on SCT dataset 43.85 47.23
The TTA process introduces an additional 0.23 seconds per image, with a performance improvement of +3.38 PSNR. This trade-off is typical for TTA methods and demonstrates effective adaptation to unseen data. The overhead arises from the model updating its weights to better align with the target domain distribution, without requiring ground-truth labels. This extra computation is common in test-time adaptation methods (Kundu et al., 2020; Liu et al., 2023a; Wang et al., 2022; Sun et al., 2020; Wang et al., 2020), as they rely on on-the-fly optimization to adapt to distribution shifts during inference.
Table 18: Metrics to assess the diversity of different HDR datasets.
FHLP Fraction of HighLight Pixel (Guo et al., 2023)
EHL Extent of HighLight (Guo et al., 2023)
SI Spatial Information (Series, 2012)
CF ColorFulness (Hasler and Suesstrunk, 2003)
stdL standard deviation of Luminance (Guo et al., 2023)
ALL Average Luminance Level (Guo et al., 2023)
DR Dynamic Range: the log10 differences between the highest 2% luminance and the lowest 2% luminance. (Hu et al., 2022b)